OpenStudy (anononion):

Quick question about factoring. ​f(x)=2x^3 − 3x^2 −5x +​6; k=1. K=1 is a factor. So I synthetically divide to get the new polynomial: 2x^2 - x- 6 My question is silly. When I set up to factor how do I know the order. (x )(2x ) or (2x )(x ) and more, if my terms are -3 and 2 how do I know which goes where without testing it all with FOIL

OpenStudy (simontheengineer):

You can use a quadratic equation to find out your two values? This is usually what is done to find out two values for x.

OpenStudy (anononion):

well normally I look at the last number and find the factors that will add to the middle term, my question is how do I know which is written first, a -3 or 2 for example.

OpenStudy (mathmale):

The order in which you write your factors is immaterial.

OpenStudy (mathmale):

You are correct: the quotient is 2x^2 - x- 6. You can now apply synth. div. to finding the roots of 2x^2 - x- 6. Write your 3 factors as (x-1)(x-?)(x+?). You could check your work by multiplying out this expression.

OpenStudy (watchmath):

If you want you can do like this: First write \(2x^2-x-6\) as \(\frac{1}{2}\left((2x)^2-2x-12\right)\). Let \(y=2x\), so you want to factor \(\frac{1}{2} (y^2-y-12)=\frac{1}{2}(y+3)(y-4)=(2x+3)(x-2)\).